# GCR: Geometry-Consistent Routing for Task-Agnostic Continual Anomaly Detection

### v1

@@ 1 Introduction
Feature-based anomaly detection (AD) has become a widely adopted paradigm in industrial visual inspection, owing to its strong empirical performance and practical accessibility: a pre-trained vision encoder provides transferable patch representations, and anomalies are detected by measuring deviations from normal feature patterns without requiring task-specific supervision [bergmann2019mvtec, defard2021padim, roth2022towards]. A key implication of this paradigm is that, when the feature space is sufficiently informative, competitive detection and localization can be achieved without end-to-end representation learning, using simple scoring rules on top of frozen embeddings [defard2021padim, roth2022towards].
However, real industrial systems rarely operate in a single-category regime. A single production line often manufactures dozens of products whose visual patterns differ subtly, and test-time inference is frequently performed without access to category labels or reliable meta-information. This motivates task-agnostic anomaly detection, where a system must handle multiple normal manifolds while making decisions without category supervision at inference time [you2022unified, gao2024learning, he2024diffusion, he2024learning].
In addition, product portfolios evolve: new categories are introduced over time, and industrial inspection systems are expected to incorporate new normal patterns while maintaining performance on previously deployed categories. This motivates continual anomaly detection, where categories arrive sequentially and the system must avoid severe degradation on earlier categories after learning new ones [liu2024unsupervised]. A recurring empirical observation in such settings is that introducing new categories can dramatically reduce performance on past categories under task-agnostic evaluation, a phenomenon often described as catastrophic forgetting at the system level [liu2024unsupervised]. Recent approaches such as DFM and UCAD address this challenge by adopting multi-head designs and routing mechanisms, and further attempt to improve routing via additional representation learning or auxiliary objectives [wu2025dfm, liu2024unsupervised].
This raises a fundamental question: if the feature space is already sufficiently discriminative, is additional representation learning truly necessary for routing stability? In a multi-head system, once per-category experts are constructed, overall success under task-agnostic continual evaluation depends critically on expert selection—routing each test input to an appropriate head before applying any head-specific anomaly scoring [liu2024unsupervised]. Therefore, it is natural to ask whether the observed continual collapse primarily reflects insufficient representational capacity, or whether it can largely be attributed to decision instability in cross-head routing when head-wise scores are heterogeneous [you2022unified, liu2024unsupervised].
To probe this hypothesis, we conduct a diagnostic study using PatchCore-style retrieval on top of frozen patch embeddings from a pre-trained ViT backbone. We construct per-category prototype banks from normal training patches and perform task-agnostic inference by routing each test image through a shared geometric criterion in the frozen embedding space, followed by standard within-head anomaly scoring [roth2022towards]. Surprisingly, this routing strategy yields near-no performance degradation relative to category-aware evaluation in our setup, suggesting that the frozen feature space provides sufficient separability for both category discrimination and anomaly detection when routing is geometry-consistent [roth2022towards]. This observation, together with the strong sensitivity of continual performance to routing behavior reported in prior work, suggests that a substantial portion of the “catastrophic forgetting” observed in task-agnostic continual AD may be better understood as cross-head decision instability rather than representational degradation [liu2024unsupervised].
@@
Motivated by this perspective, we propose GCR (Geometry-Consistent Routing) for task-agnostic continual anomaly detection. GCR adopts a multi-head Mixture-of-Experts view and decouples cross-head routing from within-head anomaly scoring: routing is performed using a shared geometric criterion in the frozen patch-embedding space, and anomaly maps are computed only within the selected expert. By focusing on routing design as a decision-rule problem, GCR improves routing stability under heterogeneous heads without requiring end-to-end representation learning [liu2024unsupervised, you2022unified].
@@
  • A decision-rule perspective on task-agnostic continual AD. We view task-agnostic continual anomaly detection as a multi-head expert selection problem and discuss how performance collapse under continual evaluation can be associated with cross-head decision instability in routing, in addition to representational factors [you2022unified, liu2024unsupervised].
  • GCR: geometry-consistent routing as a lightweight MoE design. We propose GCR, a lightweight Mixture-of-Experts framework that separates cross-head routing from within-head anomaly scoring. GCR routes inputs using a shared geometric criterion in a frozen patch-embedding space with per-category prototype banks, and evaluates anomaly scores only within the selected expert, avoiding direct cross-head comparisons of heterogeneous head scores [roth2022towards, liu2024unsupervised].
  • Evaluation with routing diagnostics. We evaluate GCR on MVTec AD and VisA under task-agnostic continual protocols and report routing diagnostics (e.g., routing accuracy and conditional performance) to help disentangle routing errors from within-head scoring behavior. The results indicate that geometry-consistent routing is associated with improved routing stability and reduced continual performance collapse while retaining competitive detection and localization performance [bergmann2019mvtec, liu2024unsupervised, wu2025dfm].
@@ 2.1 Industrial Anomaly Detection
Industrial Anomaly Detection (IAD) aims to detect and localize visual defects under the constraint that only normal samples are available during training. This setting reflects real-world manufacturing scenarios and is typically addressed in unsupervised or weakly supervised regimes. Since the introduction of the MVTec AD benchmark, most studies have focused not only on image-level anomaly classification but also on pixel-level defect localization, where the representation of normal data and the design of anomaly scoring functions play a critical role.[bergmann2019mvtec] As the field evolves, the focus has expanded to more complex challenges through diverse datasets: VisA introduces scenarios with multiple instances and complex structures to overcome the limitations of well-aligned object views[zou2022spot], while MVTec LOCO-AD shifts the paradigm toward ”logical anomalies” where global context, such as component counts or spatial relationships, must be understood beyond simple local textures.[bergmann2022beyond] Furthermore, the MVTec 3D-AD dataset incorporates depth information to address defects identifiable only in three-dimensional space[bergmann2021mvtec], and the large-scale Real-IAD benchmark provides a more robust evaluation under varying illumination and viewing angles, pushing the boundaries of IAD toward more practical and scalable industrial applications[wang2024real].
@@
Existing approaches can be broadly categorized by how normality is defined and modeled. Early works predominantly adopted reconstruction-based methods, where anomalies are identified through reconstruction errors produced by autoencoders or related generative models. Representative examples include autoencoder variants enhanced with structural similarity measures, self-supervised segmentation networks, and discriminatively trained reconstruction embeddings. While conceptually intuitive, these methods often struggle to preserve fine-grained textures and may generalize poorly to complex industrial patterns.[bergmann2018improving, song2021anoseg, zavrtanik2021draem]Moreover, anomaly scores in these approaches are typically defined through reconstruction errors or likelihood-related quantities, which are used directly without explicit normalization across samples or evaluation settings.
@@
Subsequently, probabilistic density estimation approaches emerged, explicitly modeling the distribution of normal data in feature space.[rudolph2021same, gudovskiy2022cflow, yu2021fastflow, zhang2023unsupervised] Normalizing-flow-based methods estimate likelihoods for anomaly scoring and have demonstrated strong performance with pixel-level localization, benefiting from exact density estimation and invertible mappings. However, their training complexity and sensitivity to architectural choices remain non-trivial challenges. More recently, diffusion-based models have been explored for anomaly detection by reconstructing normal patterns or removing defects through iterative denoising processes.[wyatt2022anoddpm, zhang2025diffusionad] Extensions incorporating text conditions or prompt-driven generation further expand this line of research toward controllable anomaly synthesis and augmentation.[sun2025unseen]Although these methods provide principled probabilistic formulations, the resulting anomaly scores are often applied directly, with limited consideration of their comparability across different categories or heterogeneous feature manifolds.
@@
In parallel, feature-based anomaly detection leveraging powerful pre-trained vision encoders has become the dominant paradigm in industrial inspection. Instead of reconstructing inputs, these methods distinguish normal and anomalous samples directly in feature space by modeling the distribution or structural relationships of normal features. This category includes teacher–student knowledge distillation frameworks[wang2021student, deng2022anomaly, gu2023remembering], one-class classification methods,[zhou2021vae, yi2020patch, liu2023simplenet] explicit feature distribution modeling approaches, and memory-bank-based nearest-neighbor methods.[roth2022towards, bae2023pni] Owing to their strong performance and relatively simple pipelines, such approaches have inspired a wide range of follow-up studies.However, anomaly scores in these methods are typically defined as unnormalized distances or similarity measures in feature space, implicitly assuming that such scores remain meaningful across different local neighborhoods or categories.
@@
Recently, there has been growing interest in extending unsupervised feature-based anomaly detection to unified multi-class and continual learning settings, where multiple categories are handled jointly or introduced sequentially.[you2022unified, guo2025dinomaly] At the same time, vision–language models and large foundation models have been adopted to enable zero-shot or prompt-based anomaly detection, significantly broadening the scope of industrial inspection beyond fixed-category assumptions.[jeong2023winclip, zhou2023anomalyclip, li2024promptad, gu2024anomalygpt] These trends indicate a shift toward more flexible and scalable anomaly detection systems.While these works emphasize system-level scalability and robustness, most operate on top of predefined anomaly scoring functions, without explicitly revisiting how anomaly scores themselves are defined or normalized under heterogeneous conditions.
@@ 4.1 Experimental Setup
We evaluate GCR on MVTec AD, which contains 15 categories (objects and textures). Following the standard protocol, only normal images are used for training, while the test set contains both normal and anomalous samples. We report both image-level anomaly detection and pixel-level anomaly localization.
Feature extractor.
We use a frozen ViT-B/16 from OpenCLIP/CLIP [schuhmann2021laion]. For each input image, we extract multi-layer patch token features by recording intermediate outputs at a fixed set of transformer blocks and concatenating them along the channel dimension. The encoder is kept frozen in all experiments; all routing and anomaly scoring are performed directly in the resulting patch embedding space.
@@
We compare GCR with representative anomaly detection methods, including probabilistic modeling approaches (PaDiM [defard2021padim]), memory-based retrieval methods (PatchCore [roth2022towards], CFA [lee2022cfa]), and recent unified or continual anomaly detection frameworks (SimpleNet [liu2023simplenet], UniAD [you2022unified], UCAD [liu2024unsupervised]). For methods originally designed for category-aware inference, we follow their official implementations and evaluation protocols. Unless otherwise stated, all baselines are evaluated under the same backbone, input resolution, and dataset splits. For continual experiments, we follow the standard sequential protocol and report forgetting using image-level AUROC.
@@
For each category $c$, we construct a prototype bank $\mathcal{M}_{c}=\{\mu_{c,k}\}_{k=1}^{K}$ from normal training images. We first collect all patch features from the training set of category $c$, and then select $K$ representative prototypes using greedy $k$-center (farthest-first) coreset selection under squared Euclidean distance. Unless otherwise stated, prototypes are fixed after selection and are not updated by learning.
@@
At test time, the category identity is assumed unknown. We perform task-agnostic routing directly in the patch embedding space. For each category $c$, we compute a routing distance by accumulating patch-wise nearest-prototype distances:
@@
We route each test sample to the top-1 category $\hat{c}(x)=\arg\min_{c}r_{c}(x)$ by default. Given the routed head, we compute a patch anomaly map $S_{\hat{c}}(x)$ using our prototype-based scoring rule (LSE energy by default, with hard-min as a comparable variant). The anomaly map is upsampled to the input resolution by bilinear interpolation, and the image-level score is computed by top-$q$ pooling over pixels. Optionally, one may fuse the top-$k$ routed heads by an elementwise maximum; we treat this as a robustness option rather than the default.
@@
We use a single set of hyperparameters across all categories. The default settings in our main experiments are: topk=1 for routing and topq=0.01 for image-level pooling. For patch scoring, we use a smooth LogSumExp (LSE) aggregation over prototypes by default; hard-min nearest-prototype scoring yields comparable rank-based metrics in our setting and is included in ablations. Unless explicitly varied, we use uniform component weights $\pi_{c,k}=1/K$ and keep all computations in the frozen patch embedding space.
@@
All experiments are run on a single NVIDIA RTX 4090 GPU. Prototype selection is performed once per category offline. At inference, the dominant computation is the encoder forward pass and the evaluation of prototype energies for the routed candidate heads.
@@ 4.2 Evaluation Metrics
We adopt the standard forgetting measure (FM) in continual learning [chaudhry2018riemannian]. We consider a sequential setting where categories arrive in order $t=1,\dots,T$. At step $t$, we build a new head (prototype bank and associated scoring parameters) for the new category and keep previously constructed heads unchanged. After each step $t$, we evaluate the system on each previously seen category $i\leq t$ using the task-agnostic inference pipeline (i.e., routing among the currently available heads; optional top-$k$ fusion is reported separately when used), and compute a performance score $P_{i}^{(t)}$. The forgetting for category $i$ is
@@
Class CFA [lee2022cfa] PaDiM [defard2021padim] PatchCore [roth2022towards] SimpleNet [liu2023simplenet] UniAD [you2022unified] UCAD [liu2024unsupervised] GCR (ours)
@@
Class CFA [lee2022cfa] PaDiM [defard2021padim] PatchCore [roth2022towards] SimpleNet [liu2023simplenet] UniAD [you2022unified] UCAD [liu2024unsupervised] GCR (ours)
@@
Class CFA [lee2022cfa] PatchCore[roth2022towards] SimpleNet [liu2023simplenet] UniAD [you2022unified] UCAD [liu2024unsupervised] GCR (ours)
@@
Class CFA [lee2022cfa] PatchCore [roth2022towards] SimpleNet [liu2023simplenet] UniAD [you2022unified] UCAD [liu2024unsupervised] GCR (ours)
@@ 4.3 Ablation Study
Ablation I: Routing rule is the dominant factor.
We first ablate the routing decision rule while keeping the backbone frozen and using identical per-category prototype banks. We compare: (a) score-based routing, which selects the head by comparing head-specific anomaly scores (e.g., $\arg\min_{c}\,\mathrm{score}_{c}(x)$), and (b) geometry-consistent routing, which selects the head by accumulated nearest-prototype distances in the patch embedding space,
where $\mathcal{C}_{t}$ is the set of available heads at continual step $t$ and $N$ is the number of patch tokens.

### v2

@@ 1 Introduction
Feature-based anomaly detection (AD) is widely used in industrial inspection because a pre-trained vision encoder yields transferable patch embeddings, enabling detection and localization via simple deviation scoring without task-specific supervision or end-to-end training [3, 7, 21].
In practice, however, inspection systems must handle multiple product categories and often operate without category labels at inference time, i.e., under a task-agnostic setting [30, 9, 14, 15]. Moreover, product categories arrive sequentially, requiring continual incorporation of new normal patterns while preserving performance on previously deployed ones. Under task-agnostic continual evaluation, prior work reports severe performance drops on earlier categories, commonly described as system-level catastrophic forgetting [19].
We argue that a major driver of this collapse can be cross-head decision instability rather than representational degradation. In multi-head systems, overall success hinges on expert selection (routing) before applying any head-specific anomaly scoring; heterogeneous head scores can make this selection unstable even with a fixed encoder [19, 30]. To test this, we conduct a diagnostic study using PatchCore-style retrieval on frozen ViT patch embeddings. By routing test inputs with a shared geometry-consistent criterion in the embedding space and performing anomaly scoring only within the selected head, we observe near-zero degradation relative to category-aware evaluation in our setup [21].
@@
Motivated by this observation, we propose GCR (Geometry-Consistent Routing) for task-agnostic continual AD. GCR adopts a Mixture-of-Experts view and explicitly decouples cross-head routing from within-head anomaly scoring: routing is performed via a shared geometric criterion in the frozen patch-embedding space, and anomaly maps are computed only within the selected expert. This design improves routing stability under heterogeneous heads without additional representation learning.
@@
  • A decision-rule perspective on task-agnostic continual AD. We view task-agnostic continual anomaly detection as a multi-head expert selection problem and discuss how performance collapse under continual evaluation can be associated with cross-head decision instability in routing, in addition to representational factors.
  • GCR: geometry-consistent routing as a lightweight MoE design. We propose GCR, a lightweight Mixture-of-Experts framework that separates cross-head routing from within-head anomaly scoring. GCR routes inputs using a shared geometric criterion in a frozen patch-embedding space with per-category prototype banks, and evaluates anomaly scores only within the selected expert, avoiding direct cross-head comparisons of heterogeneous head scores.
  • Evaluation with routing diagnostics. We evaluate GCR on MVTec AD and VisA under task-agnostic continual protocols and report routing diagnostics (e.g., routing accuracy and conditional performance) to help disentangle routing errors from within-head scoring behavior. The results indicate that geometry-consistent routing is associated with improved routing stability and reduced continual performance collapse while retaining competitive detection and localization performance.
@@ 2.1 Industrial Anomaly Detection
Industrial Anomaly Detection (IAD) aims to detect and localize visual defects under the constraint that only normal samples are available during training. This setting reflects real-world manufacturing scenarios and is typically addressed in unsupervised or weakly supervised regimes. Since the introduction of the MVTec AD benchmark, most studies have focused not only on image-level anomaly classification but also on pixel-level defect localization, where the representation of normal data and the design of anomaly scoring functions play a critical role.[3] As the field evolves, the focus has expanded to more complex challenges through diverse datasets: VisA introduces scenarios with multiple instances and complex structures to overcome the limitations of well-aligned object views[37], while MVTec LOCO-AD shifts the paradigm toward ”logical anomalies” where global context, such as component counts or spatial relationships, must be understood beyond simple local textures.[2] Furthermore, the MVTec 3D-AD dataset incorporates depth information to address defects identifiable only in three-dimensional space[4], and the large-scale Real-IAD benchmark provides a more robust evaluation under varying illumination and viewing angles, pushing the boundaries of IAD toward more practical and scalable industrial applications[26].
@@
Existing approaches can be broadly categorized by how normality is defined and modeled. Early works predominantly adopted reconstruction-based methods, where anomalies are identified through reconstruction errors produced by autoencoders or related generative models. Representative examples include autoencoder variants enhanced with structural similarity measures, self-supervised segmentation networks, and discriminatively trained reconstruction embeddings. While conceptually intuitive, these methods often struggle to preserve fine-grained textures and may generalize poorly to complex industrial patterns.[5, 24, 32]Moreover, anomaly scores in these approaches are typically defined through reconstruction errors or likelihood-related quantities, which are used directly without explicit normalization across samples or evaluation settings.
@@
Subsequently, probabilistic density estimation approaches emerged, explicitly modeling the distribution of normal data in feature space.[22, 12, 31, 34] Normalizing-flow-based methods estimate likelihoods for anomaly scoring and have demonstrated strong performance with pixel-level localization, benefiting from exact density estimation and invertible mappings. However, their training complexity and sensitivity to architectural choices remain non-trivial challenges. More recently, diffusion-based models have been explored for anomaly detection by reconstructing normal patterns or removing defects through iterative denoising processes.[28, 33] Extensions incorporating text conditions or prompt-driven generation further expand this line of research toward controllable anomaly synthesis and augmentation.[25]Although these methods provide principled probabilistic formulations, the resulting anomaly scores are often applied directly, with limited consideration of their comparability across different categories or heterogeneous feature manifolds.
@@
In parallel, feature-based anomaly detection leveraging powerful pre-trained vision encoders has become the dominant paradigm in industrial inspection. Instead of reconstructing inputs, these methods distinguish normal and anomalous samples directly in feature space by modeling the distribution or structural relationships of normal features. This category includes teacher–student knowledge distillation frameworks[27, 8, 11], one-class classification methods,[36, 29, 20] explicit feature distribution modeling approaches, and memory-bank-based nearest-neighbor methods.[21, 1] Owing to their strong performance and relatively simple pipelines, such approaches have inspired a wide range of follow-up studies.However, anomaly scores in these methods are typically defined as unnormalized distances or similarity measures in feature space, implicitly assuming that such scores remain meaningful across different local neighborhoods or categories.
@@
Recently, there has been growing interest in extending unsupervised feature-based anomaly detection to unified multi-class and continual learning settings, where multiple categories are handled jointly or introduced sequentially.[30, 13] At the same time, vision–language models and large foundation models have been adopted to enable zero-shot or prompt-based anomaly detection, significantly broadening the scope of industrial inspection beyond fixed-category assumptions.[16, 35, 18, 10] These trends indicate a shift toward more flexible and scalable anomaly detection systems.While these works emphasize system-level scalability and robustness, most operate on top of predefined anomaly scoring functions, without explicitly revisiting how anomaly scores themselves are defined or normalized under heterogeneous conditions.
@@ 4.1 Experimental Setup
We evaluate GCR on two standard anomaly detection benchmarks: MVTec AD and VisA. MVTec AD consists of 15 categories covering both objects and textures, while VisA contains 12 industrial object categories. For both datasets, only normal images are used for training, and the test sets contain both normal and anomalous samples. We report image-level anomaly detection and pixel-level anomaly localization results following the standard evaluation protocols of each benchmark.
Feature extractor (ours).
Unless otherwise stated, our method uses a frozen OpenCLIP/CLIP ViT-B/16 pretrained on LAION-400M (ViT-B/16, laion400m_e32) [23] as the feature extractor. For an input image, we record patch-token features from a fixed set of transformer blocks and concatenate them along the channel dimension to form the patch embedding. The encoder is kept frozen in all experiments; all routing and anomaly scoring are performed directly in this fixed patch embedding space.
@@
We compare GCR with representative anomaly detection methods, including probabilistic modeling approaches (PaDiM [7]), memory-based retrieval methods (PatchCore [21], CFA [17]), and recent unified or continual anomaly detection frameworks (SimpleNet [20], UniAD [30], UCAD [19]). For all baselines, we follow their official implementations and default evaluation protocols. Backbone architectures, input resolutions, and training configurations may differ across methods due to their original design choices; unless explicitly stated, we do not claim backbone-matched or fully controlled comparisons.
@@
For each category $c$, we construct a prototype bank $\mathcal{M}_{c}=\{\mu_{c,k}\}_{k=1}^{K}$ using only normal training images. We collect all patch features from the training set of category $c$ and select $K$ representative prototypes via greedy $k$-center (farthest-first) coreset selection under squared Euclidean distance. Unless otherwise specified, the selected prototypes are fixed after construction and are not updated by learning.
@@
At test time, the category identity of a sample is assumed unknown. We perform task-agnostic routing directly in the patch embedding space. For each category $c$, we compute a routing distance by accumulating patch-wise nearest-prototype distances:
@@
Each test sample is routed to the top-1 category $\hat{c}(x)=\arg\min_{c}r_{c}(x)$ by default. Given the routed category, we compute a patch-level anomaly map $S_{\hat{c}}(x)$ using our prototype-based scoring rule. We use a smooth LogSumExp (LSE) energy aggregation over prototypes by default, and include hard-min nearest-prototype scoring as a comparable variant. The anomaly map is upsampled to the input resolution by bilinear interpolation, and the image-level anomaly score is obtained via top-$q$ pooling over pixels. Optionally, multiple routed heads can be fused by an elementwise maximum, which we treat as a robustness variant rather than the default setting.
@@
We use a single set of hyperparameters across all categories and both datasets. Unless otherwise stated, the default settings are topk=1 for routing and topq=0.01 for image-level pooling. For patch scoring, we use LogSumExp aggregation with uniform component weights $\pi_{c,k}=1/K$.
@@
All experiments for GCR are conducted on a single NVIDIA RTX 4090 GPU. Prototype selection is performed once per category offline. At inference time, the dominant computation consists of a forward pass through the frozen encoder and prototype energy evaluation for the routed candidate heads.
@@ 4.2 Evaluation Metrics
We adopt the standard forgetting measure (FM) in continual learning [6]. We consider a sequential setting where categories arrive in order $t=1,\dots,T$. At step $t$, we build a new head (prototype bank and associated scoring parameters) for the new category and keep previously constructed heads unchanged. After each step $t$, we evaluate the system on each previously seen category $i\leq t$ using the task-agnostic inference pipeline (i.e., routing among the currently available heads; optional top-$k$ fusion is reported separately when used), and compute a performance score $P_{i}^{(t)}$. The forgetting for category $i$ is
@@
Class CFA [17] PaDiM [7] PatchCore [21] SimpleNet [20] UniAD [30] UCAD [19] GCR (ours)
@@
Class CFA [17] PaDiM [7] PatchCore [21] SimpleNet [20] UniAD [30] UCAD [19] GCR (ours)
@@
Class CFA [17] PatchCore[21] SimpleNet [20] UniAD [30] UCAD [19] GCR (ours)
@@
Class CFA [17] PatchCore [21] SimpleNet [20] UniAD [30] UCAD [19] GCR (ours)